I posted a question previously but worded it wrong.
It was supposed to be find the values of the six trigonometric functions of an angle O(the o with the line through it) in standard posistion whos terminal side passes through point P
with P being (-7,24)

Answer by stanbon(48558)

find the values of the six trigonometric functions of an angle O(the o with the line through it) in standard posistion whos terminal side passes through point P, with P being (-7,24).
------------
Plot that point on an x/y coordinate system; notice it is in the 2nd Quadrant.
Draw a vertical line segment from the point to the x-axis.
Dras a segment connecting P to the origin.
You have a right triangle with x = -7, y = 24
Then r = sqrt((-7)^2+(24)^2) = 25
----------------
sin(theta) = y/r = 24/25 ; csc(theta) = 25/24
cos(theta) = x/4 = -7/25 ; sec(theta) = -25/7
tan(theta) = 24/-7 ; ctn(theta) = -7/24
=================================================
Cheers,
Stan H.